OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP Lynn Wessel borrowed $2,500 for 18 months. The total interest she paid was $315. What rate of interest did Lynn pay?

OpenStudy (doc.brown):

Do you want the rate to be in dollars per month?

OpenStudy (doc.brown):

Or probably percent per month?

OpenStudy (anonymous):

Yes looking for the percentage rate per month

OpenStudy (doc.brown):

So how much was paid per month?

OpenStudy (anonymous):

Ya if we can figure out the interest percentage for one month then Im guessing we can multiple it by the 18 months

OpenStudy (doc.brown):

Or if you take the total paid and divide it by 18 months it will tell you dollar amount paid in one month.

OpenStudy (anonymous):

Then how do I find the Percent?

OpenStudy (anonymous):

This seems simple Interest. Formula is Interest = Principal Amount x Rate x Time. Rate and Time must relate. Since you're taking time to be in months, your rate will come out to be monthly too.

OpenStudy (anonymous):

however we don't know the rate ?

OpenStudy (anonymous):

Comes out to be 17.50/month but then how do I find the percentage rate?

OpenStudy (doc.brown):

The rate in 18 months is \(\frac{$315}{18mo}\) so the rate in one month is \(\frac{$17.50}{mo}\). What percent is $17.50 of $2500?

OpenStudy (anonymous):

Rate will come out as Percentage per month.

OpenStudy (anonymous):

70% ?

OpenStudy (anonymous):

I= PRT 315=2500(R)(18) Simplify for Rate, it will be percentage per month

OpenStudy (doc.brown):

If you have $2500 and I steal 70% of it, did I steal $17.50?

OpenStudy (anonymous):

no

OpenStudy (doc.brown):

I think about fractions like this, how many pieces make a whole goes on the bottom, and the number of pieces I have go on top. In this case the whole is $2500 and the pieces Lynn had to pay was $17.50.\[\frac{17.50}{2500}\]When you divide it you're reducing the whole to 1.\[\frac{17.50}{2500}=\frac{0.007}{1}\]But in the case of percent you want the whole to be 100 pieces big.\[\frac{0.007}{1}\times\frac{100}{100}=\frac{0.7}{100}\]Or 0.7 per cent (cent meaning 100).

OpenStudy (anonymous):

Ok that helps thank you!

OpenStudy (doc.brown):

You can do it the other way too,\[\frac{315}{2500}\times 100=12.6\%\]That's the total interest in percent.\[\frac{12.6\%}{18mo}=\frac{0.7\%}{mo}\]